# Denis Roio Dyne.org &Rebecca Selvaggini

### v1

@@ Abstract
It is of critical importance to design digital identity systems that ensure the privacy of citizens as well protect them from issuer corruption. Unfortunately, what Europe’s and USA’s public sectors are currently developing does not offer such basic protections. Aiming at solving this issue, we propose a method for untraceable selective disclosure and privacy preserving revocation of digital credentials, utilizing the unique homomorphic characteristics of second order Elliptic Curves and Boneh-Lynn-Shacham (BLS) signatures operated on them. Our approach ensures that users can selectively reveal only the necessary credentials, while protecting their privacy across multiple presentations. We also aim to protect users from issuer corruption, by making it possible to apply a threshold for revocation to require collective agreement among multiple revocation issuers.
@@ 1 Introduction
Such systems are known as selective disclosure and this article aims at improving their cryptography to adhere to basic privacy-by-design standards.
@@ 1.1 State of the art
Selective disclosures are being used by nation states across the world in their next generation identity systems, for instance EIDAS2.0 in Europe where the European Digital Identity Wallet Architecture and Reference Framework[1] mandates the use of SD-JWT[2] and mDOC [3]. Unfortunately both SD-JWT and mDOC adopt for their cryptography just Hash-Based Message Authentication Codes (HMAC) to generate proofs, and they can be easily traced (non-unlinkability): recent critical feedback on the EUDI ARF[4] details several problems.
In North America the efforts concentrate on the adoption of the the BBS+ algorithm[5] leveraging its Zero Knowledge Proof properties and applied to W3C Verifiable Credentials[6] to obtain an higher degree of privacy.
@@ 1.2 Threats considered
The EUDI-ARF standard dictates that credentials issued to a holder, can be presented (in the form of a verifiable presentation) to a relying party in order to have one or more attributes verified. Every verifiable presentation includes one or more HMAC(s), formatted in SD-JWT: the HMACs are identical each time a verifiable presentation is produced from a certain credential. This makes possible for colluding relying parties, or to malicious actors, to trace a holder’s identity by collecting, exchanging and confronting verifiable presentations. In order to preserve the privacy of a credential holder the disclosures presented by the wallet should not be traceable across different presentations (unlinkability). This threat appears to be well mitigated by BBS+ through its Zero Knowledge Proof implementation.
@@
There is no privacy-preserving revocation system designed, either in EUDI-ARF, or in W3C-VC and BBS+. The approach followed for the revocation is that of the status (revocation) lists [7] for W3C-VC and EUDI-ARF is investigating a similar method. The use of a certificate status lists (CRL) presents issues related primarily to privacy [8], because sensitive information about holders leaks from the list. This is partially mitigated by expiration dates, in cases where credentials can be short-lived (typically less important credentials), but not applicable with digital identification documents such as ID, driving license, passport and social security numbers, which typically have longer or no expiration time. In case the choice of strategy for revocation is left open to developers, the risk for major privacy breaches may occur, for example with the adoption of public status lists [9]. Based on these arguments, we believe that anonymous revocations are a sine qua none condition to guarantee a sufficient level of privacy in digital identities and credentials. Our approach is that of designing a privacy-preserving revocation mechanism to remove the leak of holder’s information.
@@
If the choice of interactive revocation is left to a single issuer, one may unilaterally choose to revoke credentials, without being subject to revision or having to seek consensus with a quorum of issuers. This situation leads to security issues in case Issuers are corrupted and make a weaponized use of digital revocations to persecute engaged individuals. Such a condition becomes a real concern for journalists or activists living under dictatorial regimes that may arbitrarily revoke their credentials, or even ID cards and passports. Similarly, a security breach of an issuer service, would result in similar threats. We mitigate this risk by introducing the possibility for threshold issuance of revocation keys.
@@ 2.1 Key contributions
The cryptographic scheme described in the paper, name SD-BLS for brevity, takes inspiration by modern zero knowledge proofs schemes such as [5]BBS signatures or Coconut[10]. The paper aims to employ some of the properties of the said schemes, overcome some of the shortcomings of SD-JWT and mDOC formats, mainly the non-unlinkability of credentials, proposing a novel cryptographic approach to similar data structures. Furthermore SD-BLS proposes novel anonymous cryptographic revocation flow for verifiable credentials, that aims to solve the privacy issues posed by status and revocation lists.
Unlinkabe selective disclosure
Similarly to the SD-JWT and mDOC formats, SD-BLS produces an array of claims: the elements of the array are individually signed by the issuer. In SD-BLS the signature(s) replace the HMAC. Using the homomorphic properties of this curve SD-BLS enables the holder to produce a different presentation each time they want to prove a claim. SD-BLS can be used to prove credentials in public (on blockchain) when they are boolean attributes.
@@
SD-BLS proposes a novel approach to credential revocation: the data published by the revocation issuer will produce cryptographic material that contains no information about and is never linkable to the holder. The cryptographic revocation material allows anyone to verify if an SD-BLS proof produced by a credential holder has been revoked by one or multiple issuers. The unlinkability, and thus anonymity of the cryptographic revocation, allows the revocation issuer to publish revocations off or on blockchain.
@@ 2.2 Applications
In this section we present some applications and use cases for digital identity and credentials, that could benefit from using the unlinkability and anonymous revocation of the the SD-BLS scheme.
@@
The focus of the EUDI-ARF specifications is identity documents: it defines mechanisms and data structures to issue a Personal Identification (PID) as well as digital driving licenses. Similarly, the US government is experimenting with W3C-VC and mDOC for cross-states interoperable driving licenses. SD-BLS data format is similar to SD-JWT and mDOC, offering selective disclosure and adding unlinkability and anonymous revocation.
@@
We are unaware of standardization efforts for interoperable credentials in the fields of "Know Your Customer" (KYC) and Anti Money-Laundering (AML) certifications. We are aware of solution providers experimenting with W3C-VC for AML applications.
@@
SD-BLS can be used on blockchain to produce publicly verified proofs when the credentials are made of clear-text messages stating a boolean property or ACL clearance.
  • A verifiable presentation can be used by a smart-contract to verify if one or more holder’s claim match the requirement needed to process a transaction. The verifiable presentation is anonymous, therefore its storage in the transaction log doesn’t pose privacy issues and is GDPR compatible.
  • Issuers can publish their cryptographic revocation lists on chain, allowing smart-contracts to verify the status of a credential, still granting unlinkability.
@@
It’s worth mentioning that, while the verifiable presentation is itself unlinkable to the holder’s blockchain address, the unlinkability of the holder to their blockchain address at verification phase, needs to be addressed within the blockchain protocol used: this issue is out of scope in this paper.
@@ 3.2 Issuance
We create a new secret key $rev$ with public key $r_{2}$, then compute the point $r_{1}\in\mathbb{G}_{1}$ and signature of the claim $\sigma_{rev}$ to add to the issuer signature:
$\begin{split}r_{2}&=rev\cdot G_{2}\\ r_{1}&=rev\cdot G_{1}\\ \sigma_{rev}&=rev\cdot U\\ \sigma&=sk\cdot U+\sigma_{rev}\\ \end{split}$
@@
$\mathcal{C}=\big{\{}\{r_{1i},r_{2i},\sigma_{i}\}:i\in claims\big{\}}$
@@ 3.3 Presentation
As we have done in the issuance phase (3.2), we exploit the homomorphic property of BLS signatures once again, but this time to render the presented signature unlinkable (impossible to trace across multiple presentations) by adding a "blinding factor" signature to the one provided by the issuer.
Let $\mathcal{H}$ denote a cryptographic hash function (e.g. SHA-256) and let $\varepsilon$ be an integer chosen uniformly at random that serves as the blinding factor. Using the same technique of the Tripartite Diffie-Hellman Key Exchange [12] we construct a secret that is binded to the blinding factor, the Issuer public key and the "revocation point" $r_{1}=rev\cdot G_{1}$ computed in the issuance phase:
$\tau=\mathcal{H}(e(A.pk,r_{1})^{\varepsilon})$ (1)
When the holder needs to disclose a particular claim $m$ we construct the credential proof as follow: let $\sigma$ be the signature of the claim provided by the Issuer, then apply the blinding factor to the signature using the secret $\tau$
$s=\sigma+sign(\tau,m)$ (2)
and the corresponding public key computed as
$p=r_{2}+G_{2}\cdot\tau$
where $r_{2}$ is the revocation key linked to the claim. We also attach the public key $r=G_{1}\cdot\varepsilon$ that correspond to the secret key $\varepsilon$.
@@ 3.4 Verification
Credential verification is made by checking that the issuer has signed the blinded proof being presented. In order to do that the credential issuer’s public key must be added to the proof.
@@
$c_{p}=\{id,p,r,s\}$
where $s$ and $p$ are respectively the signature and the public key of the credential proof of the string $id$ after verifying the status of the revocation, we can check the validity of the presented claim computing the key
$pk=A.pk+p$
and verify the bls signature.
@@ 3.4.1 Proof of correctness
The signature $s$ is given by
$\begin{split}s&=\sigma+sign(\tau,id)\quad\text{where}\\ \sigma&=sign(A.sk,id)+sign(rev,id)\\ \text{so}\quad s&=A.skU+revU+\tau U\end{split}$
@@
$\begin{split}e(pk,U)&=e(A.skG_{2}+revG_{2}+\tau G_{2},U)\\ &=e(A.skG_{2}+revG_{2})\cdot e(\tau G_{2},U)\\ &=e(A.skG_{2},U)\cdot e(revG_{2},U)\cdot e(\tau G_{2},U)\\ &=e(G_{2},A.skU)\cdot e(G_{2},revU)\cdot e(G_{2},\tau U)\\ &=e(G_{2},A.skU)\cdot e(G_{2},revU+\tau U)\\ &=e(G_{2},A.skU+revU+\tau U)\\ &=e(G_{2},s)\\ \end{split}$
@@ 3.5 Revocation
To make sure no revocation has been emitted for each credential, we update the revocation list from the issuer. A revocation list can be publicly distributed since revocation keys do not provide any information on the identity of holders, unless they are presented in a verification.
Given an element $c_{p}=\{id,p,r,s\}$ of the credential proof table and the Issuer public key $A.pk$, if the claim id is present in the revocation list we can take the corresponding revocation private key $rev$ and compute the factor $\tau$ as
$\tau=\mathcal{H}(e(A.pk,r)^{rev})$ (3)
The equation (3) gives the same result as (1) since using the bilinearity of the pairing [12] we have:
$\begin{split}\tau&=\mathcal{H}(e(A.pk,r)^{rev})\\ &=\mathcal{H}(e(A.pk,G_{1}\cdot\varepsilon)^{rev})\\ &=\mathcal{H}(e(A.pk,G_{1})^{rev\cdot\varepsilon})\\ &=\mathcal{H}(e(A.pk,G_{1}\cdot rev)^{\varepsilon})\\ &=\mathcal{H}(e(A.pk,r_{1})^{\varepsilon})\end{split}$
Then we can verify if a credential claim has been revoked checking first if the revocation public key match:
$G_{2}\cdot rev=p-G_{2}\cdot\tau$
and then verify the following signature with the issuer’s public key (pk):
$sig=s-sign(\tau,id)-sign(rev,id)$ (4)
This signature should be verified with the issuer’s pk since:
[…]
then exploiting the homomorphic property of BLS signature we got that $sig$ in (4) is signed using the issuer’s private key.
@@ 3.6 Threshold Revocation
In order to split responsibilities over interactive revocation we introduce an optional threshold over the revocation key, plus we split the functionality of credential issuance from that of revocation issuance, now operated by different peers.
This is implemented via an interactive process facilitated by a third party trustee who should be:
@@
Such a trustee will be in contact with credential issuers for the signature of credentials: it will create the $rev$ revocation key while concealing it from them. The trustee then proceeds creating the $\sigma_{rev}=rev\cdot U$ revocation signature and the $r_{1}=rev\cdot G_{1}$ projection as detailed in the issuance phase (3.2). Finally the revocation trustee will distribute $\sigma_{rev}$ and $r_{1}$ to issuers for further processing into the final signed credential.
The trustee will then proceed asynchronously to split the revocation key into shares using a public verifiable secret sharing (PVSS [13]) implementation and distribute these shares to all revocation issuers. In order to issue a revocation, a configurable quorum of peers among the revocation issuers will need to reconstruct the secret revocation key and publish it.
@@ 4 Benchmarks
Lab measurements on a growing number of claims show that issuance is less computation heavy and that the production of proofs and their verification need approximately the same resources, as shown in figure 1.
Figure 1: Speed comparison of issue, prove and verify
Based on our benchmarks, the resulting data objects sizes are:
  • Claim: 322 Bytes
  • Proof: 226 Bytes
  • Revocation: 64 Bytes
The computational cost grows for the verification phase with the presence of a cryptographic revocations list. We assume that revocations are grouped by claim type or credential structure, for instance all revocations of a certain claim (e.g.: "is above 18" or "is a Danish citizen") for any holder are published in the same list: this will not degrade the level of privacy and will group all the revocations in smaller lists.
@@
At last we measure the binary hamming distance between unlinkable proofs of the same credential signed by the same issuer. We performed this test without using any special random generator. This measurement should serve as a quick evaluation of the "randomness" of all blinded contents constituting a proof: the signature, the issuer’s public key and the revocation’s public key, from this benchmark the claim’s identifier is excluded. The results are shown in figure 3 and include an overlay of frequency values taken from the comparison of two random values. This benchmarks show that the binary hamming distance between different proofs is the same as the one between two random values of the same length generated by the same PRNG.
Figure 3: Frequency of hamming distance between unlinkable proofs
@@ 5.1 Security considerations
The revocations database is privacy and corruption sensitive (by our previous definition of issuer corruption) and it should be securely stored by each Issuer. This is mitigated by the optional adoption of threshold revocation. When using threshold revocation there is still the need for a revocation trustee who may be a single point of failure if allowed to publish revocation keys dishonestly collected during the process of credential issuance.
BLS signatures and the proof system obtained with credentials are considered secure by assuming the existence of random oracles [14], together with the decisional Diffie-Hellman Problem (DDH) [15], the external Diffie-Hellman Problem (XDH), and with the Lysyanskaya-Rivest-Sahai-Wol Problem (LRSW) [16], which are connected to the Discrete Logarithm. In fact, under these assumptions, we have that our protocol satisfies unforgeability, blindness, and unlinkability.

### v2

@@ Abstract
It is of critical importance to design digital identity systems that ensure the privacy of citizens as well as protecting them from issuer corruption. We aim to solve this issue and propose a method for selective disclosure and privacy preserving revocation of digital credentials, using the unique homomorphic characteristics of second order Elliptic Curves and Boneh-Lynn-Shacham (BLS) signatures. Our approach ensures that users can selectively reveal credentials signed by a certain issuer, which can be interactively revoked by a quorum of other agreeing issuers without revealing the identity of users. Our goal is to protect users from issuer corruption by requiring collective agreement among multiple revocation issuers.
@@ 1 Introduction
Such systems are known as selective disclosure and they enhance users privacy by allowing data minimization when proving credentials.
@@ 1.1 State of the art
Selective disclosures are being used by nation states across the world in their next generation identity systems, for instance EIDAS2.0 in Europe where the European Digital Identity Wallet Architecture and Reference Framework[1] mandates the use of SD-JWT[2] and mDOC [3]. SD-JWT adopts for its cryptography Hash-Based Message Authentication Codes (HMAC) to generate proofs: such presentations can be traced and recent critical feedback on the EUDI ARF[4] details this problem.
In North America the efforts concentrate on the adoption of the BBS+ algorithm[5] leveraging its Zero Knowledge Proof properties and applied to W3C Verifiable Credentials[6] to obtain an higher degree of privacy by making every disclosure unlinkable.
@@ 1.2 Threats considered
The EUDI-ARF standard dictates that credentials issued to a holder, can be presented (in the form of a verifiable presentation) to a relying party in order to have one or more attributes verified. Every verifiable presentation includes one or more HMAC(s), formatted in SD-JWT: the HMACs are identical each time a verifiable presentation is produced from a certain credential. This makes possible for colluding relying parties, or to malicious actors, to trace a holder’s identity by collecting, exchanging and confronting verifiable presentations (linkability). This threat appears to be well mitigated by BBS+ through its Zero Knowledge Proof implementation.
@@
We believe that anonymous revocations are a condicio sine qua non to guarantee a sufficient level of privacy in digital identities and credentials. There is no privacy-preserving revocation system designed, either in EUDI-ARF, or in W3C-VC and BBS+. In case the choice of strategy for revocation is left open to developers, the risk for major privacy breaches may occur, for example with the adoption of public status lists [7]. The hypotetical use of a certificate status lists (CRL) presents issues related primarily to privacy [8], because sensitive information about holders leaks from the list. This problem is partially mitigated by expiration dates, in cases where credentials can be short-lived (typically less important credentials), but not applicable with digital identification documents such as ID, driving license, passport and social security numbers, which typically have longer or no expiration time. Future plans for the national standards we are observing include the adoption of “Bitstring” status lists [9] which may grant a degree of privacy. Our approach is that of designing a privacy-preserving revocation mechanism to remove the leak of holder’s information and allow to delegate the governance of revocations to multiple revocation issuers which may be different from the credential issuer.
@@
If the choice of interactive revocation is left to a single issuer, one may unilaterally choose to revoke credentials, without being subject to revision or having to seek consensus with a quorum of issuers. This situation leads to security issues in case Issuers are corrupted and make a weaponized use of digital revocations to persecute engaged individuals. Such a condition becomes a real concern for journalists or activists living under dictatorial regimes that may arbitrarily revoke their credentials, or even ID cards and passports. Similarly, a security breach of an issuer service, would result in similar threats. We mitigate this risk by introducing the possibility for threshold issuance of revocation keys and by separating the responsibility of revocation issuance and credential issuance.
Feature Comparison
UP UR TR
SD-BLS no yes yes
SD-JWT no wip no
BBS+ yes wip no
Table 1: feature comparison
We briefly round up on feature differences between the named selective disclosure cryptographic schemes, as shown in table 1, mainly distinguishing between three fundamental features:
  • UP: Unlinkable Presentation
  • UR: Unlinkable Revocation
  • TR: Threshold Revocation
Where wip is mentioned meaning work in progress on adoption of bitstring status lists for unlinkable revocations.
@@ 2.1 Key contributions
The cryptographic scheme described in this paper, named SD-BLS for brevity, implements some of the properties of the SD-JWT scheme (which can be easily encoded in mDOC format) and proposes a novel cryptographic approach to similar data structures. Furthermore SD-BLS proposes novel anonymous cryptographic revocation flow for verifiable credentials, that aims to solve the privacy issues posed by status and revocation lists.
Selective disclosure
Similarly to the SD-JWT and mDOC formats, SD-BLS produces an array of claims: the elements of the array are individually signed by the issuer. In SD-BLS the signature(s) replace the HMAC. Using the homomorphic properties of this curve SD-BLS enables the holder to selectively disclose only certain issuer signed claims and produce a presentation that minimizes private information given to verifiers.
@@
SD-BLS proposes a novel approach to credential revocation: the data published by the revocation issuer will produce cryptographic material that contains no information about the credential holders. The cryptographic revocation material allows anyone to verify if an SD-BLS proof produced by a credential holder has been revoked by one or multiple issuers. The unlinkability, and thus anonymity of the cryptographic revocation, allows the revocation issuer to share revocations in public and allows anyone to verify if credentials have been revoked.
@@ 2.2 Applications
In this section we present some applications and use cases for digital identity and credentials, that could benefit from using the SD-BLS scheme.
@@
The focus of the EUDI-ARF specifications is identity documents: it defines mechanisms and data structures to issue a Personal Identification (PID) as well as digital driving licenses. Similarly, the US government is experimenting with W3C-VC and mDOC for cross-states interoperable driving licenses. SD-BLS data format is similar to SD-JWT and mDOC, offering selective disclosure and anonymous revocation.
@@
We are unaware of standardization efforts for interoperable credentials in the fields of "Know Your Customer" (KYC) and Anti Money-Laundering (AML) certifications. We are aware of solution providers experimenting with W3C-VC for AML applications and believe SD-BLS can greatly improve the governance of credential revocation, which is a critical component for this use case.
@@
SD-BLS can be used with blockchain-based smart-contracts to activate certain functions:
  • A verifiable presentation can be used by a smart-contract to verify if one or more holder’s claim match the requirement needed to process a transaction.
  • Issuers can publish their cryptographic revocation lists on chain, allowing smart-contracts to verify the status of a credential.
@@ 3.2 Issuance
We create a new secret key $rev$ with public key $r$ and signature of the claim $\sigma_{rev}$ to add to the issuer signature:
$\begin{split}r&=rev\cdot G_{2}\\ \sigma_{rev}&=rev\cdot U\\ \sigma&=sk\cdot U+\sigma_{rev}\\ \end{split}$
@@
$\mathcal{C}=\big{\{}\{id,r,\sigma\}:id\in claims\big{\}}$
@@ 3.3 Presentation
When the holder needs to disclose a particular claim $id$ it is sufficient to send the signed claim set $\{id,r,\sigma\}$.
@@ 3.4 Verification
Credential verification is made by checking the presented issuer’s signature. In order to do that the credential issuer’s public key must be added to the revocation public key.
@@
$c_{id}=\{id,r,\sigma\}$
where $\sigma$ and $r$ are respectively the signature of the string $id$ and the revocation public key. We can check the validity of the presented claim computing the key
$pk=A.pk+r$
and verify the bls signature $\sigma$.
@@ 3.4.1 Proof of correctness
The signature $\sigma$ is given by
$\begin{split}\sigma&=sign(A.sk,id)+sign(rev,id)\\ &=A.skU+revU\end{split}$
@@
$\begin{split}e(pk,U)&=e(A.skG_{2}+revG_{2},U)\\ &=e(A.skG_{2},U)\cdot e(revG_{2},U)\\ &=e(G_{2},A.skU)\cdot e(G_{2},revU)\\ &=e(G_{2},A.skU+revU)\\ &=e(G_{2},\sigma)\\ \end{split}$
@@
This concludes the first verification phase. If the given presentation is valid, then the verifier should proceed to check the revocation status of the credential as explained below.
@@ 3.5 Revocation
To control if a revocation has been emitted for any credential being verified, we update the revocation list from the issuer. A revocation list can be publicly distributed since revocation keys do not provide any information on the identity of holders.
Given an element $c_{id}=\{id,r,\sigma\}$ of the credential presentation and the Issuer public key $A.pk$, if the claim id is present in the revocation list, we can take the corresponding revocation private key $rev$.
We can verify the validity of the revocation by checking if the revocation public key presented by the holder match:
$r=G_{2}\cdot rev$
This step should follow the signature verification and is sufficient to guarantee the revocation status.
If a dishonest holder provides a wrong $r$, then the verification of the signature will fail, and the given credential should be considered invalid.
If it happens that the signature verification is successful and that the revocation key does not match, one can conclude that the credential is not revoked.
@@ 3.6 Threshold Revocation
In order to split responsibilities over interactive revocation we introduce a threshold over the revocation key, plus we split the functionality of credential issuance from that of revocation issuance, now operated by different peers.
This is implemented via an interactive process facilitated by a third party trustee, a revocation dealer, who should be:
@@
Such a revocation dealer will be in contact with credential issuers for the signature of credentials: it will create the $rev$ revocation key while concealing it from them. The trustee then proceeds creating the $\sigma_{rev}=rev\cdot U$ revocation signature and the $r=rev\cdot G_{2}$ public key as detailed in the issuance phase (3.2). Finally the revocation trustee will distribute $\sigma_{rev}$ and $r$ to issuers for further processing into the final signed credential.
The revocation trustee (dealer) will then proceed asynchronously to split the revocation key into shares using a public verifiable secret sharing (PVSS [11]) implementation and distribute these shares to all revocation issuers. In order to issue a revocation, a configurable quorum of peers among the revocation issuers will need to reconstruct the secret revocation key and publish it.
This process separates responsibilities between the credential issuer and the revocation issuers, delegating to the revocation issuers the possibility to revoke a credential interactively through a collective process.
The collection of shares can be done asynchronously and is provable. The dealer should publish proofs of knowledge of each revocation share, proving their creation and authenticity. Revocation issuers can also use the dealer proofs to verify the validity of the shares received without revealing them. Such revocation proofs will also be useful when revocation issuers will reconstruct a revocation, since they can refer to them to prove their shares are authentic without revealing their content.
@@ 4 Benchmarks
Lab measurements on a growing number of claims show that issuance is less computation heavy than verification, as shown in figure 1.
Figure 1: Speed comparison of issue and verify
Based on our benchmarks, the resulting data objects sizes, excluding the signed material which can be either of variable length or fixed length hashes, are:
  • Proof: 145 Bytes
  • Revocation: 32 Bytes
The computational cost grows for the verification phase with the presence of a cryptographic revocations list. We assume that revocations are grouped by claim type or credential structure, for instance all revocations of a certain common claim (e.g.: "is above 18" or "is a Danish citizen") for any holder are published in the same list: this will not degrade the level of privacy and will group some revocations in smaller lists.
@@
The threshold operated for revocation issuers consists of a verifiable secret sharing implementation supporting a configurable total and quorum of peers. Our implementation shows very good performance on reconstruction, which is the most speed-sensitive operation in scenarios where responsive revocation process is required. The process of reconstruction can be easily scaled for asynchronous consensus on a micro-service swarm architecture and verified on blockchain.
Figure 3: Speed of creation and reconstruction of shares among multiple peers
@@ 5.1 Security considerations
The revocations database is privacy and corruption sensitive (by our previous definition of issuer corruption) and it should be securely stored by each Issuer. This is mitigated by the adoption of threshold revocation. When using threshold revocation there is still the need for a revocation trustee who may be a single point of failure if allowed to publish revocation keys dishonestly collected during the process of credential issuance.
BLS signatures and the proof system obtained with credentials are considered secure by assuming the existence of random oracles [12], together with the decisional Diffie-Hellman Problem (DDH) [13], the external Diffie-Hellman Problem (XDH), and with the Lysyanskaya-Rivest-Sahai-Wol Problem (LRSW) [14], which are connected to the Discrete Logarithm.
@@
The SD-BLS implementation we present in this paper doesn’t integrate any protection against replay-attacks, nor implements timestamps and expiration. We believe security features protecting from replay-attacks can be easily added and adapted from the context of deployment. I absence of session management providing such protections, any real-world use of SD-BLS should contemplate adding a signed timestamp to presentations and expiration checks on verifications.